# Through the Perspective of LiDAR: A Feature-Enriched and Uncertainty-Aware Annotation Pipeline for Terrestrial Point Cloud Segmentation

### v1

@@ 2.2.3 Stage 3: Back Projection and Refinement in 3D Space
Figure 8: Illustration of the stage 3 processes.
@@
Stage 3 connects the 2D annotation workflow back to the 3D domain (Fig. 8). Each 2D segmentation mask $L(i,j)$ is back-projected to its corresponding 3D point coordinates, producing an labeled point cloud. Fig. 9 illustrates examples of colorized 3D renderings from different feature groups, enabling visual inspection and cross-validation. A key challenge arises from class ambiguity along object boundaries in the 2D projection (e.g., stem–canopy transitions). After back-projection, however, each point is re-embedded within its geometric neighborhood, providing local context for disambiguation. To leverage this, we apply a two-stage refinement process in 3D space: 1) Geometric smoothing via $k$-nearest neighbor (kNN) majority voting to suppress small boundary errors; 2) Feature-driven repair using a Random Forest classifier trained on a reliable core set, which corrects systematic ambiguities along complex boundaries. Together, these steps yield crisp and reliable 3D semantic annotations, which we further verify through manual inspection.
@@
Technical details of the back-projection equations, kNN voting, and Random-Forest relabeling are provided in D(Eqs. 1015), along with additional information and applications of the virtual spheres.
@@ 2.3 Performance Evaluation and Analysis
We evaluate each feature set and its epistemic-uncertainty map with the metrics in Table 2. Segmentation quality is captured by overall accuracy (oAcc), mean class accuracy (mAcc), mean IoU (mIoU) and per-class IoU (IoUc). Pixel-wise errors and their uncertainty are quantified with Shannon Entropy, while the uncertainty map’s ability to expose errors is measured by the area under the precision–recall curve (AUPRC) between the uncertainty map and a binary error mask.
@@
To assess the data efficiency of our annotation pipeline, we perform experiments with progressively larger subsets of annotated training data. Specifically, we randomly sample training sets of size $n=\{4,\,8,\,12,\,16,\,20,\,24,\,28\}$ scans. For each subset size, we maintain a fixed validation ratio of 0.25, ensuring integral train/validation set splits (e.g., 3/1 for $n=4$; 6/2 for $n=8$, etc.). The test set remains 9 scans from plots #6–7 across all experiments.
@@ 3 Results
Fig. 11 illustrates several consistent patterns across the four representative test scans. It is clear that the ensemble model captures vertical stratification well: the upper canopy (green), root zone (orange), and ground-and-water layer (purple) predictions are almost indistinguishable from the ground-truth masks, indicating that the model handles the dominant classes with high confidence.
Errors concentrate along cluttered, high-frequency boundaries, particularly in two transition zones: (i) the interface between the distant upper canopy and the fine stems that pierce it, where sparse sampling and rapid color–depth changes reduce discriminative power; and (ii) the contact between exposed roots and the muddy, debris-strewn ground, where heterogeneous textures dominate. Most of these challenging pixels are captured by the epistemic-uncertainty maps. Bright-red patches in the uncertainty layer generally coincide with the white error pixels beneath, yielding AUPRC scores of 0.30–0.40. The initial segment of each precision–recall curve starts near unity, showing that the top 3–5 % most-uncertain pixels are predominantly true errors—valuable for targeted manual inspection in an active-learning workflow. Despite this alignment, prediction reliability gaps remain. As highlighted by purple rectangles, discrepancies are most pronounced in scans with dense root tangles (e.g., site07), where AUPRC drops from 0.403 to 0.302.
@@ 3.2 Effect of feature enrichment
Across single-channel inputs—raw intensity, range, Z, their preprocessed versions, and per-point geometry (curvature, anisotropy, planarity)—mean IoU ranges between 0.702 - 0.731. Most three-channel combinations perform similarly, but the contrast-enhanced I.R.Z stack (intensity, range, inverse Z) lifts mIoU to 0.745, confirming the value of the histogram stretch in preprocessing. Merging feature groups into a six-channel input pushes performance further (0.754–0.761), yet adding still more channels yields no real gain: results level off at $\approx 0.76$. The best configuration, IRZ_N3_CAP, tops out at 0.768—three points better than raw intensity alone and six points better than raw range. In practice, therefore, a compact six-channel feature set captures nearly all available discriminative power on the Mangrove3D dataset, whereas larger stacks add computation complexity with little return.
Examining performance per class, Ground & Water and Void classes consistently achieve high IoU scores ($\geq 0.80$) across nearly all configurations, demonstrating their ease of separation irrespective of feature complexity. Conversely, the Stem class remains consistently challenging (IoU ranging between 0.47 - 0.53), highlighting persistent segmentation difficulties, due to fewer sampling points as distance increases and its thin, partially occluded structures. The Canopy, Roots, and Objects classes notably benefit from enriched feature sets, with improvements of 0.07–0.10 IoU points observed when moving from single- to six-channel inputs. The best overall balance is achieved with the six-channel IRZ_PCA, providing the highest IoU for four of the six semantic classes and achieving the top mIoU (0.766).
@@
As shown in Table 3, the XYZ-only baseline achieves 0.634 mIoU, reflecting common StemCanopy confusion. Adding geometric descriptors—surface normals and C.A.P (six additional channels)—boosts performance to 0.712 mIoU, the best configuration observed. Class-level trends mirror this pattern: Objects show the largest gain (0.516 to 0.837), while Ground is already strong and improves only marginally. Beyond this six-extra-channel setup, additional features do not yield further benefits, indicating feature saturation once orientation and local-shape cues are included.
@@ 3.4.1 ForestSemantic
The ForestSemantic dataset comprises $\sim\!720$ million points sampled from six nominal 32 m $\times$ 32 m plots in Evo, Finland (61.19° N, 25.11° E). Each plot is labeled into six classes: Ground, Trunk, First-order branch, Higher-order branch, Foliage, and Miscellany. At the time of writing (Sept 2025) only three plots—#1, #3, and #5—were publicly available on Zenodo mspace_lab_2025_15193973 ; our experiments therefore focused on these.
@@
As shown in Table 4, the comparison of feature groups on the boreal-forest ForestSemantic plots reproduces the additive pattern. Replacing the raw intensity–range–height triplet with its preprocessed counterpart (I.R.Z) yields only a marginal lift over the baseline (mIoU 0.455 vs 0.453), indicating that simple radiometric normalization alone is insufficient for these cluttered, canopy-rich plots. Pure geometry (C.A.P) is far more valuable, adding +2.6 pts to mIoU (0.479) and +3.3 pts to mean class accuracy. Fusing normals with radiometry (N3) helps trunk discrimination, but does not surpass the geometric stack globally. The turning point again comes from feature fusion: adding curvature to radiometry (IRZ_CAP) pushes mIoU to 0.496, while the nine-channel union (IRZ_N3_CAP) breaks the 0.50 barrier (mIoU = 0.501). Appending a three-component PCA projection (IRZ_N3_CAP_PCA) delivers the overall best scores (oAcc = 0.902, mAcc = 0.595, mIoU = 0.511), confirming that modest dimensionality reduction can still help when the scene is dominated by fine-scale, self-occluding vegetation.
Multi-channel stacks excel at woody material: IRZ_CAP attains the best Foliage IoU (0.679) and, together with N3_CAP, boosts Higher-order branch discrimination beyond 0.015—triple the value achieved by geometry or radiometry alone. Normal information remains critical for Trunk (0.640 with IRZ_N3_CAP_PCA) and Ground (0.833). Interestingly, the PCA-extended model improves Foliage again to 0.693 and edges out competitors in seven of the eight valid classes, while maintaining the strongest overall accuracy. Taken together, the results show that in dense forest plots the geometric features (C.A.P and normals) are indispensable, and that a compact PCA head can reconcile these diverse cues without diluting their class-specific strengths.
@@ 3.4.2 Semantic3D
The Semantic3D dataset contains nearly 1.96 billion points of eight classes - 1: man-made terrain, 2: natural terrain, 3: high vegetation, 4: low vegetation, 5: buildings, 6: hardscape, 7: scanning artifacts, 8: cars. In our experiments, only the training set was used, as the ground-truth labels for the original test set are not publicly available. We split the 15 training scenes into training, validation, and test subsets in an 11:1:3 ratio.
Since the Semantic3D has an apparent long-tail class imbalance, we applied a two-stage hybrid subsampling approach. This approach first applies stratified sampling that preferentially samples rare classes ($\leq 1\%$ frequency) with a 5% allocation from the 12M target points, preventing loss of minority classes like scanning artifacts and cars that are critical for urban scene understanding. Subsequent voxel grid downsampling (0.01 m) ensures spatial uniformity while maintaining the enhanced class distribution, enabling scalable processing of large urban point clouds for semantic segmentation applications.
@@
Figure 19 renders the 3D point clouds colored by the original RGB values and the proposed feature groups. Similar to the observations on 2D feature maps, the proposed feature maps provide additional insights into object characteristics compared to original RGB or Intensity alone, such as ground vs. non-ground objects (I.R.Z stack), edges (C.A.P stack), orientations (Normals stack), and combinations of these enhancements (PCA stack).
Figure 20 visualizes virtual 3D “rings” back-projected from the colorized feature maps - with the top and bottom cropped ($Zenith\in[45^{\circ},136^{\circ}]$) from what used to be “spheres”.
@@
Table 5: Comparison of feature–input combinations for semantic segmentation on Semantic3D. Within each metric column, the top three scores are highlighted with progressively darker blue shading (dark = best, medium = second, light = third).
@@
3 baseline-RGB 0.790 0.519 0.414 0.911 0.616 0.282 0.769 0.091 0.653 0.116 0.000 0.286 0.352
3 irz_raw 0.784 0.495 0.388 0.918 0.594 0.322 0.650 0.054 0.644 0.173 0.003 0.091 0.317
3 I.R.Z 0.818 0.551 0.451 0.914 0.651 0.455 0.740 0.019 0.672 0.131 0.044 0.432 0.393
3 C.A.P 0.814 0.521 0.406 0.918 0.680 0.441 0.786 0.066 0.614 0.128 0.016 0.342 0.348
3 N3 0.893 0.646 0.562 0.927 0.823 0.767 0.837 0.181 0.675 0.183 0.047 0.612 0.516
3 PCA 0.848 0.622 0.499 0.917 0.739 0.557 0.828 0.090 0.666 0.145 0.033 0.515 0.447
3 MNF 0.810 0.547 0.424 0.891 0.672 0.486 0.758 0.138 0.571 0.179 0.041 0.084 0.369
3 ICA 0.822 0.551 0.440 0.872 0.717 0.502 0.772 0.146 0.629 0.113 0.017 0.170 0.386
6 IRZ_CAP 0.870 0.584 0.493 0.938 0.780 0.680 0.793 0.077 0.626 0.104 0.061 0.382 0.438
6 IRZ_N3 0.888 0.611 0.501 0.932 0.823 0.760 0.825 0.208 0.703 0.187 0.043 0.032 0.448
6 N3_CAP 0.877 0.615 0.507 0.934 0.798 0.698 0.814 0.207 0.669 0.100 0.048 0.292 0.453
9 IRZ_N3_CAP 0.898 0.660 0.563 0.948 0.826 0.759 0.830 0.197 0.692 0.262 0.065 0.487 0.515
12 IRZ_N3_CAP_PCA 0.877 0.614 0.512 0.939 0.787 0.701 0.822 0.178 0.635 0.166 0.043 0.333 0.458
Feature codes: baseline-RGB = original RGB values; irz_raw = intensity, range, $Z$ (raw); I.R.Z = pre-processed intensity, range, $Z_{\text{inv}}$; C.A.P = curvature, anisotropy, planarity; N3 = pseudo-RGB from surface normals; PCA = first three principal components of PCA; MNF = first three components of MNF; ICA = first three components of ICA; IRZ_CAP = I.R.Z + C.A.P; IRZ_N3 = I.R.Z + N3; N3_CAP = N3 + C.A.P; IRZ_N3_CAP = I.R.Z + N3 + C.A.P; IRZ_N3_CAP_PCA = I.R.Z_N3_cap + PCA of that set.
As summarized in Table 5, the comparative study on the urban Semantic3D benchmark echoes—and strengthens—the trends we observed in forest scenes.
Under-matched training conditions, replacing the RGB baseline (mIoU = 0.414) with the preprocessed stack of I.R.Z - which is essentially depth-aware radiometry, already lifts mean IoU by +3.7 pts and mean class accuracy by +3.2 pts, showing that well-normalized LiDAR channels can serve as a practical substitute when cameras are absent or their colors are unreliable. Pure geometry features - C.A.P - deliver a slightly lower accuracy (mIoU = 0.406). Surface-normal colorization is most impactful: the three-channel N3 stack jumps to mIoU = 0.562—an absolute +14.8 pts over RGB—while posting the second-best overall accuracy (0.893). Fusing all three sources (IRZ_N3_CAP, nine channels) yields column-best scores in every global metric (mIoU = 0.563, oAcc = 0.898, mAcc = 0.660). The statistical feature groups - p3, m3, and c3 - showed various improvements compared with baseline, +8.5pts, +1 pt, and +2.6 pts, respectively. Adding a PCA projection (IRZ_N3_CAP_PCA) offers no further benefit, indicating that the nine-channel fusion already captures the complementary information.
Normals excel on organic shapes: N3 alone tops High vegetation (IoU = 0.837) and Cars & trucks (0.612). Geometric stacks aid fine, man-made structures: IRZ_N3 pushes Buildings to IoU = 0.703. The nine-channel fusion wins or ties in seven of nine classes and more than triples IoU for Scanning artifacts class (0.065 vs. 0.017–0.048 elsewhere). Together, these results demonstrate that our feature-enrichment strategy scales beyond mangrove scenes, providing consistent, cross-domain improvements, while preserving the pipeline’s low-annotation ethos.

### v3

@@ 2.2.3 Stage 3: Back Projection and Refinement in 3D Space
(a) Workflow overview of stage 3.
(b) Before refinement.
(c) After refinement.
(d) Label histograms.
Figure 9: Stage-3 back-projection and 3D label refinement. (a) Overall workflow connecting 2D spherical annotations to the 3D point cloud via back-projection and re-projection. (b-c) Example point clouds before and after 3D refinement, showing reduced isolated label noise and improved boundary consistency. (d) Class-wise label distributions before and after refinement, indicating that local corrections are achieved while the global class balance is preserved.
@@
Stage 3 connects the 2D annotation workflow back to the 3D domain (Fig. 9). While spherical projection enables efficient processing, it may introduce localized ambiguities when multiple points map to the same pixel, particularly along object boundaries such as stem-canopy transitions. To compensate for this loss of explicit 3D neighborhood context, we apply a two-step refinement in 3D space. First, geometric smoothing is performed using k-nearest neighbor (kNN) majority voting, which suppresses small, isolated label errors by enforcing local spatial consistency. Second, a Random Forest classifier trained on reliable core regions is used to correct systematic boundary ambiguities using geometric and radiometric features. Because both stages operate directly in 3D space, they re-embed each point within its local neighborhood and mitigate projection-induced inconsistencies. Together, these steps yield crisp and reliable 3D semantic annotations, which we further verify through manual inspection. Fig. 10 illustrates examples of colorized 3D renderings from different feature groups, enabling visual inspection and cross-validation.
@@
Technical details of the back-projection equations, kNN voting, and Random-Forest relabeling are provided in D(Eqs. 10-15), along with additional information and applications of the virtual spheres.
@@ 2.3 Performance Evaluation and Analysis
We evaluate each feature set and its epistemic-uncertainty map with the metrics in Table 4. Segmentation quality is captured by overall accuracy (oAcc), mean class accuracy (mAcc), mean IoU (mIoU) and per-class IoU (IoUc). Pixel-wise errors and their uncertainty are quantified with Shannon Entropy, while the uncertainty map’s ability to expose errors is measured by the area under the precision-recall curve (AUPRC) between the uncertainty map and a binary error mask.
@@
To assess the data efficiency of our annotation pipeline, we perform experiments with progressively larger subsets of annotated training data. Specifically, we randomly sample training sets of size $n=\{4,\,8,\,12,\,16,\,20,\,24,\,28\}$ scans. For each subset size, we maintain a fixed validation ratio of 0.25, ensuring integral train/validation set splits (e.g., 3/1 for $n=4$; 6/2 for $n=8$, etc.). The test set remains 9 scans from plots #6-7 across all experiments.
@@ 3 Results
Fig. 12 shows that segmentation errors are primarily concentrated in cluttered transition regions, such as canopy–stem interfaces and root–ground boundaries. These error-prone areas are consistently highlighted by elevated epistemic uncertainty, with bright uncertainty responses largely overlapping misclassified pixels and yielding AUPRC values of 0.30–0.40 across scans. Notably, the highest-uncertainty pixels achieve near-unity precision at low recall, indicating that a small fraction of uncertain regions captures a large proportion of true errors, which is well suited for targeted human inspection in uncertainty-guided annotation workflows.
@@ 3.2 Effect of feature enrichment
Across single-channel inputs—raw intensity, range, Z, their preprocessed versions, and per-point geometry (curvature, anisotropy, planarity)—mean IoU ranges between 0.702 - 0.731. Most three-channel combinations perform similarly, but the contrast-enhanced I.R.Z stack (intensity, range, inverse Z) lifts mIoU to 0.745, confirming the value of the histogram stretch in preprocessing. Merging feature groups into a six-channel input pushes performance further (0.754-0.761), yet adding still more channels yields no real gain: results level off at $\approx 0.76$. The best configuration, IRZ_N3_CAP, tops out at 0.768—three points better than raw intensity alone and six points better than raw range. In practice, therefore, a compact six-channel feature set captures nearly all available discriminative power on the Mangrove3D dataset, whereas larger stacks add computation complexity with little return.
Table 5: Computational cost and segmentation performance of evaluated models using the Intensity–Range–Z (IRZ; 3-channel) and Intensity–Range–Z with surface normals (IRZ_N3; 6-channel) feature groups.
Feature group Model Params (M) GFLOPs Train time / epoch (s) Peak GPU mem (MB) Best epoch Peak GPU mem (GB) Inference time (s) mIoU
IRZ UNet++ 26.08 53.97 16.04 7313.7 37 7.14 0.034 0.7448
DeepLabV3+ 11.68 8.85 22.77 9645.9 67 9.42 0.045 0.6937
SegFormer 13.68 10.85 15.84 4649.5 52 4.54 0.026 0.6893
Ensemble 0.082 0.7599
IRZ_N3 UNet++ 53.70 155.69 20.55 13960.0 66 13.63 0.064 0.7646
DeepLabV3+ 22.88 9.07 19.92 18420.8 55 17.99 0.055 0.7175
SegFormer 27.09 18.49 19.52 8452.4 28 8.25 0.043 0.7029
Ensemble 0.144 0.7766
Notes: GFLOPs denote the theoretical number of floating-point operations per forward pass, estimated at an input resolution of $544\times 352\times ch$. Runtime measurements use a batch size of 10. Inference time is reported per TLS scan. All models were trained and evaluated on an HPC system using a single NVIDIA A100 GPU, with 20 GB of system memory requested per SLURM job. Peak GPU memory refers to the maximum allocated memory observed during training. Best epoch corresponds to early stopping triggered with a patience of five epochs based on validation performance.
Table 5 reports model complexity, computational cost, and segmentation accuracy across feature groups. All backbone models were trained independently, and inference was performed at the level of individual TLS scans. Inference times range from 26–64 ms for single models and remain below 150 ms per scan for the ensemble, indicating that the additional cost of ensembling is modest in practical terms. While individual architectures exhibit different efficiency–accuracy tradeoffs, the ensemble provides consistently higher accuracy across feature groups, making it well suited for generating reliable pseudo-labels in offline processing pipelines.
@@
As shown in Table 6, the XYZ-only baseline achieves 0.634 mIoU, reflecting common Stem-Canopy confusion. Adding geometric descriptors—surface normals and C.A.P (six additional channels)—boosts performance to 0.712 mIoU, the best configuration observed. Class-level trends mirror this pattern: Objects show the largest gain (0.516 to 0.837), while Ground is already strong and improves only marginally. Beyond this six-extra-channel setup, additional features do not yield further benefits, indicating feature saturation once orientation and local-shape cues are included.
@@ 3.4.1 ForestSemantic
The ForestSemantic dataset comprises $\sim\!720$ million points sampled from six nominal 32 m $\times$ 32 m plots in Evo, Finland (61.19° N, 25.11° E). Each plot is labeled into six classes: Ground, Trunk, First-order branch, Higher-order branch, Foliage, and Miscellany. At the time of writing (Sept 2025) only three plots—#1, #3, and #5—were publicly available on Zenodo Lab [2025]; our experiments therefore focused on these.
@@
Table 7 shows that geometric features dominate performance on the ForestSemantic plots, while radiometric preprocessing alone yields only marginal gains. Performance improves consistently with feature fusion, with the best results achieved by the IRZ_N3_CAP_PCA configuration (mIoU = 0.511), indicating that combining complementary geometric and radiometric cues is most effective in complex forest scenes. This performance level is broadly comparable in magnitude to results reported in the dataset reference paper for state-of-the-art methods evaluated on simpler 3-class (trunk–branch–foliage) settings Liang et al. [2025].
@@ 3.4.2 Semantic3D
The Semantic3D dataset comprises approximately 1.96 billion points annotated into eight semantic classes Hackel et al. [2017]: (1) man-made terrain, (2) natural terrain, (3) high vegetation, (4) low vegetation, (5) buildings, (6) hardscape, (7) scanning artifacts, and (8) cars.
Prior to spherical projection, we applied a two-stage hybrid subsampling strategy. First, stratified sampling was used to preferentially retain rare classes (frequency $\leq 1\%$), allocating 5% of a 12M-point target budget to mitigate the loss of minority categories such as scanning artifacts and cars. Second, voxel grid downsampling with a resolution of 0.01 m was applied to enforce spatial uniformity while preserving the enriched class distribution. This design facilitates scalable processing of large-scale urban point clouds while maintaining semantic diversity. During training, 20% of the data were held out for validation, and final evaluation was conducted on all 15 test scans.
@@
Figure 20 renders the 3D point clouds colored by the original RGB values and the proposed feature groups. Similar to the observations on 2D feature maps, the proposed feature maps provide additional insights into object characteristics compared to original RGB or Intensity alone. Figure 21 visualizes virtual 3D “rings” back-projected from the colorized feature maps - with the top and bottom cropped ($Zenith\in[45^{\circ},136^{\circ}]$) from what used to be “spheres”.
@@
Table 8: Comparison of feature-input combinations for semantic segmentation on Semantic3D. Within each metric column, the top three scores are highlighted with progressively darker blue shading (dark = best, medium = second, light = third).
@@
3 RGB 0.782 0.477 0.388 0.763 0.581 0.463 0.293 0.215 0.755 0.141 0.036 0.245 0.341
3 irz_raw 0.746 0.383 0.300 0.780 0.497 0.214 0.255 0.046 0.694 0.098 0.056 0.058 0.240
3 I.R.Z 0.783 0.509 0.411 0.778 0.569 0.428 0.360 0.072 0.763 0.133 0.313 0.285 0.365
3 C.A.P 0.807 0.513 0.422 0.817 0.604 0.425 0.398 0.154 0.803 0.178 0.216 0.204 0.436
3 N3 0.799 0.545 0.451 0.764 0.597 0.509 0.440 0.168 0.799 0.139 0.300 0.342 0.412
3 PCA 0.777 0.461 0.381 0.757 0.569 0.373 0.370 0.089 0.780 0.112 0.178 0.199 0.334
3 MNF 0.730 0.457 0.351 0.719 0.471 0.356 0.316 0.191 0.709 0.106 0.168 0.119 0.305
3 ICA 0.741 0.481 0.375 0.694 0.560 0.326 0.343 0.167 0.715 0.116 0.242 0.214 0.335
6 IRZ_CAP 0.822 0.579 0.480 0.828 0.640 0.476 0.419 0.202 0.832 0.176 0.360 0.386 0.436
6 IRZ_N3 0.848 0.589 0.504 0.854 0.698 0.584 0.465 0.160 0.836 0.157 0.375 0.411 0.461
6 N3_CAP 0.832 0.576 0.490 0.813 0.629 0.539 0.584 0.217 0.841 0.199 0.233 0.357 0.450
9 IRZ_N3_CAP 0.854 0.615 0.516 0.878 0.690 0.532 0.421 0.225 0.883 0.201 0.407 0.412 0.471
12 IRZ_N3_CAP_PCA 0.836 0.585 0.496 0.830 0.647 0.527 0.467 0.178 0.857 0.206 0.357 0.390 0.454
Table 8 reports segmentation performance on the Semantic3D benchmark for different feature-input combinations. Consistent with observations in forest scenes, multi-feature configurations outperform single-feature inputs, with IRZ_N3_CAP achieving the best overall performance across global metrics. When evaluated on the official Semantic3D test set under the standard protocol, this configuration attains an mIoU of 0.516, placing it within the performance range of mid-ranked methods on the official Semantic3D benchmark website ETH Zurich [2017].